the length of a rectangle is 6 cm less than twice its width. the perimeter of the rectangle is 30 cm. what are the dimensions of the rectangle
Answer by lwsshak3(11628)

the length of a rectangle is 6 cm less than twice its width. the perimeter of the rectangle is 30 cm. what are the dimensions of the rectangle
***
let x=width of rectangle
2x-6=length of rectangle
2*(length+width)=perimeter
2(2x-6+x)=30
2(3x-6)=30
6x-12=30
6x=42
x=7
2x-6=8
width of rectangle=7 cm
length of rectangle=8 cm